The area of the triangle is 21 in,
Height is 7 in,
What is the base's length?

Answer by jim_thompson5910(35256)

A = (bh)/2

21 = (b*7)/2

21*2 = 7b

42 = 7b

42/7 = b

6 = b

b = 6


So the base is 6 inches.

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim